# FreeChunker: A Cross-Granularity Chunking Framework

### v1

@@ Abstract
Chunking strategies significantly impact the effectiveness of Retrieval-Augmented Generation (RAG) systems. Existing methods operate within fixed-granularity paradigms that rely on static boundary identification, limiting their adaptability to diverse query requirements. This paper presents FreeChunker, a Cross-Granularity Encoding Framework that fundamentally transforms the traditional chunking paradigm: the framework treats sentences as atomic units and shifts from static chunk segmentation to flexible retrieval supporting arbitrary sentence combinations. This paradigm shift not only significantly avoids the computational overhead required for semantic boundary detection, but also enhances adaptability to complex queries. Experimental evaluation on LongBench V2 demonstrates that FreeChunker achieves superior retrieval performance compared to traditional chunking methods, while significantly outperforming existing methods in computational efficiency.
Machine Learning, Natural Language Processing, Information Retrieval, Chunking
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has substantially propelled progress in natural language processing, enabling strong generalization capabilities across diverse tasks (Brown et al., 2020). However, LLMs typically rely on static, parameterized knowledge, often making them less suitable for scenarios requiring up-to-date or domain-specific information. This limitation frequently leads to hallucinated content that may deviate from factual correctness (Huang et al., 2025; Li et al., 2023).
@@
To address this limitation, RAG has emerged as a promising paradigm. By incorporating external knowledge sources, RAG enhances the generation process with more timely and verifiable content, achieving strong performance on knowledge-intensive tasks such as open-domain question answering and fact verification (Lewis et al., 2020). However, the effectiveness of RAG critically depends on the quality of retrieved content, which is substantially influenced by document chunking strategies (Ru et al., 2024). Prior research has demonstrated that chunking granularity often directly affects intra-chunk coherence, semantic coverage, and downstream generation quality. Inappropriate chunking strategies may disrupt semantic structures, potentially leading to information fragmentation or redundancy, ultimately undermining both retrieval and generation performance (Brådland et al., 2025; Lyu et al., 2025).
@@
Figure 1: Cross-Granularity Chunking FreeChunker breaks the limitation of fixed-granularity text chunking, enabling flexible extraction of sentence combinations at arbitrary granularities.
@@
To improve chunking quality, recent efforts have explored semantic-aware chunking techniques. For instance, SemanticChunker (LangChain, 2024) segments documents based on sentence embedding similarity. Meta-Chunking (Zhao et al., 2024) identifies semantic boundaries through perplexity analysis and probability-based segmentation decisions. LumberChunker (Duarte et al., 2024) leverages LLMs to predict chunk boundaries more aligned with human interpretation. However, existing approaches often lack the flexibility to adapt to diverse query intents or varying deployment constraints. Moreover, some methods rely on computationally expensive large models, where the marginal gains of semantic chunking may not always justify the computational overhead (Qu et al., 2025). More fundamentally, these approaches still operate within a single-granularity paradigm, consequently limiting their ability to simultaneously accommodate different information needs within the same document collection.
@@
To overcome the fundamental limitation of fixed chunking schemes that cannot simultaneously provide multiple granularity levels, FreeChunker is proposed in this paper—a novel Cross-Granularity Encoding Framework that introduces an alternative chunking paradigm. As illustrated in Figure 1, unlike traditional chunking methods that are typically constrained by fixed text boundaries, the proposed method enables retrievers to extract arbitrary combinations of sentences from documents, thereby providing enhanced flexibility in content retrieval. The following core contributions are offered by the proposed method:
@@
  • Computational efficiency: The framework reduces chunking overhead by treating sentences as atomic units, avoiding expensive boundary identification processes.
  • Superior performance: The method achieves competitive retrieval performance by adapting multiple granularity levels to diverse query requirements, demonstrating enhanced effectiveness across various information retrieval scenarios in LongBench V2 Task.
@@ 2 Related Work
As a foundational step in RAG, text chunking substantially influences overall system performance. The most straightforward strategy divides documents into fixed-granularity chunks, which is simple to implement but often misaligns with semantic boundaries, consequently leading to information fragmentation or redundancy (Ru et al., 2024). More critically, fixed chunking forces a single granularity choice across the entire document, thereby preventing the system from simultaneously accessing both fine-grained details and coarse-grained context that different queries may require. This granularity rigidity fundamentally limits the retrieval system’s ability to adapt to diverse information needs within a single document collection.
@@
To address the aforementioned semantic boundary issues, recent work has explored more adaptive chunking strategies. SemanticChunker (LangChain, 2024) relies on sentence-level embedding similarity to enhance intra-chunk semantic coherence by identifying natural breakpoints based on semantic similarity thresholds. Meta-Chunking (Zhao et al., 2024) proposes two adaptive segmentation algorithms leveraging large language models’ logical awareness: Perplexity Chunking identifies logical boundaries by analyzing inter-sentence perplexity variations, while Margin Sampling Chunking determines boundaries by comparing model probability differences for segmentation decisions. LumberChunker (Duarte et al., 2024) leverages large language models to determine chunk boundaries more aligned with human interpretation. However, despite their semantic awareness, these methods still operate under a single granularity paradigm—each document is ultimately represented at one chosen level of granularity, whether sentence-level, paragraph-level, or custom semantic units.
@@
Inspired by the concept of mixture of experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1994), recent methods like MoC (Zhao et al., 2025) and MoG (Mix-of-Granularity) (Zhong et al., 2025) attempt to enhance adaptability by treating different chunking strategies as expert modules and employing routing mechanisms to dynamically select the most suitable chunker based on query characteristics. However, these MoE-inspired methods essentially aggregate existing chunkers as experts; consequently, both the quality of semantic chunking and the degree of granularity control remain fundamentally bounded by the capabilities of the underlying chunker experts.
@@ 3 Methodology
In FreeChunker, sentences are treated as atomic units and chunk embeddings are generated across multiple granularities simultaneously. This approach typically enables retrieval systems to access any desired chunk size on demand, ranging from fine-grained single sentences to coarse-grained multi-sentence contexts.
@@ 3.1 Cross-Granularity Chunk Pattern
Given a document with $n$ sentences $\mathcal{D}=\{s_{1},s_{2},\dots,s_{n}\}$, a Chunk Pattern Mask $\mathbf{P}\in\{0,-\infty\}^{m\times n}$ is constructed, where each row $\mathbf{P}[i,:]$ represents the $i$-th chunk pattern, and $\mathbf{P}[i,j]=0$ typically indicates that sentence $s_{j}$ belongs to that pattern.
@@
Therefore, the framework can support any combination of chunk sizes and overlapping strategies.
@@ 3.2 Cross-Granularity Encoder
After obtaining the Chunk Pattern Mask $\mathbf{P}$, independently encoding each chunk would typically cause extensive redundant computation. Sentenizer is introduced to split documents into sentences and encode them using a token-level embedding model $\mathcal{M}$, thereby obtaining the sentence embedding matrix $\mathcal{E}=[\vec{e_{1}},\dots,\vec{e_{n}}]\in\mathbb{R}^{n\times d}$. For any chunk spanning from sentence $s_{i}$ to sentence $s_{j}$, its true embedding is denoted as $\vec{e}_{i,j}=\mathcal{M}([s_{i},s_{i+1},\dots,s_{j}])$, where $s_{i},s_{i+1},\dots,s_{j}]$ represents the concatenation of sentences.
A specialized transformer architecture is then modified. For each layer, a learnable chunk embedding $\mathbf{h}_{\text{chk}}\in\mathbb{R}^{d}$ is introduced that is replicated $m$ times to form $\mathcal{H}=[\mathbf{h}_{\text{chk}},\mathbf{h}_{\text{chk}},\dots,\mathbf{h}_{\text{chk}}]^{T}\in\mathbb{R}^{m\times d}$ and attention is used to generate $m$ chunk embeddings in parallel:
@@
The complete chunk encoding layer then applies layer normalization and feed-forward networks:
$\mathcal{V}=\text{LayerNorm}(\mathcal{H}+\text{Attn}(\mathcal{H},\mathcal{E}))$ (4)
$\text{Output}=\text{LayerNorm}(\mathcal{V}+\text{FFN}(\mathcal{V}))$ (5)
This design typically allows parallel generation of all chunk embeddings in a single forward pass, with sentence-level embeddings being reused across different granularities, thereby avoiding redundant encoding of sentences.
Figure 3: Training Dynamics Across Model Scales. Training and validation curves (top) and test set loss distributions (bottom) for three different model configurations, demonstrating consistent convergence and effective generalization.
4 Theoretical Analysis
A central concern for FreeChunker is whether the cross-granularity encoder induces non-negligible substitution loss relative to directly encoding concatenated text chunks. This section develops a geometric analysis of vector substitution, establishing worst-case and average-case guarantees that are verifiable in practice.
4.1 Setup and Constraints
Let $\vec{e}_{i,j}=\mathcal{M}([s_{i},s_{j}])$ denote the true chunk embeddings of the concatenated chunk $s_{i},s_{j}]$, and let $\vec{v}_{i,j}$ be the approximate chunk embeddings produced by the cross-granularity encoder. Let $\vec{q}$ be a query embedding. Throughout this section all vectors are assumed to be unit-norm.
The substitution loss is studied as follows:
Two parameters characterize the geometric configuration:
$\rho$ $\;=\;\cos(\vec{e}_{i,j},\vec{v}_{i,j}),$ (7)
$s$ $\;=\;\cos(\vec{q},\vec{e}_{i,j}),$
4.2 Worst-Case Substitution Bound
Theorem 1 (Vector Substitution Loss Upper Bound). Given the parameters in (7),
$\varepsilon\;\leq\;s(1-\rho)\;+\;\sqrt{1-s^{2}}\,\sqrt{1-\rho^{2}}.$ (8)
Proof. Let $\alpha=\angle(\vec{q},\vec{e}_{i,j})$ and $\beta=\angle(\vec{e}_{i,j},\vec{v}_{i,j})$, so $s=\cos\alpha$ and $\rho=\cos\beta$. By the spherical cosine law,
$\cos(\vec{q},\vec{v}_{i,j})\;=\;\cos\alpha\cos\beta+\sin\alpha\sin\beta\cos\varphi,$ (9)
where $\varphi$ is the relative azimuth of $\vec{q}$ and $\vec{v}_{i,j}$ around the axis $\vec{e}_{i,j}$. The adversarial configuration is $\cos\varphi=-1$, which yields $\cos(\vec{q},\vec{v}_{i,j})\geq s\rho-\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}$. Subtracting from $\cos(\vec{q},\vec{e}_{i,j})=s$ gives (8). $\square$
Interpretation.
The first term $s(1-\rho)$ is typically the radial component; the second term $\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}$ is tangential and becomes dominant when the query is weakly aligned with the true chunk ($s$ small) and the substitution deviates from the truth ($\rho$ small).
4.3 Expected Error under Random Azimuth
The worst-case bound (8) assumes an adversarial azimuth between $\vec{q}$ and $\vec{v}_{i,j}$. In practice, the relative azimuth $\varphi$ between these vectors around $\vec{e}_{i,j}$ is typically random rather than adversarial. A mild and empirically verifiable assumption is therefore introduced.
Assumption 1 (Random Azimuth). The relative orientation between query and approximate vectors around the true chunk embeddings is uniformly random.
Theorem 2 (Expected Substitution Loss). Under Assumption 1 and given the parameters in (7), the expected substitution loss is bounded by:
$\mathbb{E}_{\varphi}[\varepsilon\mid s,\rho]\leq s(1-\rho)+\frac{2}{\pi}\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}.$ (10)
Proof. Using $\varepsilon\leq s(1-\rho)+\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}|\cos\varphi|$ and $\mathbb{E}[|\cos\varphi|]=2/\pi$ for uniform $\varphi$ yields the bound. $\square$
Interpretation.
The bound consists of a radial term $s(1-\rho)$ and a tangential term with coefficient $2/\pi\!\approx\!0.636$, showing that random orientations substantially reduce expected degradation compared to worst-case scenarios.
Table 1: Results of different chunking methods on LongBench V2 (without Top-1 results)
Method jina-embeddings-v2-small-en nomic-embed-text-v1.5 BGE-M3
Top-5 Top-10 Time (↓) Top-5 Top-10 Time (↓) Top-5 Top-10 Time (↓)
I. Single-Document QA
Traditional 30.86% 30.86% 2.11s 35.43% 32.00% 3.76s 36.00% 34.29% 8.93s
SemanticChunker 33.71% 33.71% 12.87s 29.14% 32.57% 15.48s 32.57% 30.86% 19.83s
PPL Chunking 31.43% 28.00% 33.09s 32.57% 34.29% 34.73s 33.14% 33.71% 39.82s
Margin Sampling 30.29% 36.57% 364.64s 35.43% 36.00% 368.68s 34.29% 28.00% 375.22s
FreeChunker 32.57% 36.57% 2.75s 36.00% 36.00% 5.29s 38.29% 36.57% 9.64s
@@ 5.1 Training Setup
Three token-level embedding models are employed as encoders $\mathcal{M}$ for the Sentenizer: jina-embeddings-v2-small-en(Günther et al., 2023)(33M parameters), nomic-embed-text-v1.5 (Nussbaum et al., 2024) (137M parameters) and BGE-M3 (Chen et al., 2024) (570M parameters). Our sentence-level Cross-Granularity encoder (330M parameters) reuses partial structures and weights from BGE-M3, with dimension pruning support enabling token-level embedding model reuse across different dimensions. Documents are split into sentences and encoded to generate sentence embeddings $\mathcal{E}=[\vec{e_{1}},\dots,\vec{e_{n}}]\in\mathbb{R}^{n\times d}$.
@@
The training dataset comprises 100K documents sampled from The Pile (Gao et al., 2020). For each document, the first 8000 tokens are extracted to construct training pairs. Documents are split into sentences using the Sentenizer, typically yielding 10-200 sentences per document. Training pairs $(\vec{e},\vec{v})$ are constructed across 5 different granularities (2, 4, 8, 16, 32 sentences), where $\vec{e}$ represents ground-truth chunk embeddings obtained by directly encoding concatenated sentence spans, and $\vec{v}$ denotes the corresponding cross-granularity encoder outputs.

### v2

@@ Abstract
Chunking strategies significantly impact the effectiveness of Retrieval-Augmented Generation (RAG) systems. Existing methods operate within fixed-granularity paradigms that rely on static boundary identification, limiting their adaptability to diverse query requirements. This paper presents FreeChunker, a Cross-Granularity Encoding Framework that fundamentally transforms the traditional chunking paradigm: the framework treats sentences as atomic units and shifts from static chunk segmentation to flexible retrieval supporting arbitrary sentence combinations. This paradigm shift not only significantly avoids the computational overhead required for semantic boundary detection, but also enhances adaptability to complex queries. Experimental evaluation on LongBench V2 demonstrates that FreeChunker possesses significant advantages in both retrieval performance and time efficiency compared to existing chunking methods. The pre-trained models and codes are available at https://github.com/mazehart/FreeChunker.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has substantially propelled progress in natural language processing, enabling strong generalization capabilities across diverse tasks Brown et al. (2020). However, LLMs typically rely on static, parameterized knowledge, often making them less suitable for scenarios requiring up-to-date or domain-specific information. This limitation frequently leads to hallucinated content that may deviate from factual correctness Huang et al. (2025); Li et al. (2023).
@@
To address this limitation, RAG has emerged as a promising paradigm. By incorporating external knowledge sources, RAG enhances the generation process with more timely and verifiable content, achieving strong performance on knowledge-intensive tasks such as open-domain question answering and fact verification Lewis et al. (2020). However, the effectiveness of RAG critically depends on the quality of retrieved content, which is substantially influenced by document chunking strategies Ru et al. (2024). Prior research has demonstrated that chunking granularity often directly affects intra-chunk coherence, semantic coverage, and downstream generation quality. Inappropriate chunking strategies may disrupt semantic structures, potentially leading to information fragmentation or redundancy, ultimately undermining both retrieval and generation performance Brådland et al. (2025); Lyu et al. (2025).
@@
Figure 1: FreeChunker breaks the limitation of fixed-granularity text chunking, enabling flexible extraction of sentence combinations at arbitrary granularities.
@@
To improve chunking quality, recent efforts have explored semantic-aware chunking techniques. For instance, SemanticChunker LangChain (2024) segments documents based on sentence embedding similarity. Meta-Chunking Zhao et al. (2024) identifies semantic boundaries through perplexity analysis and probability-based segmentation decisions. LumberChunker Duarte et al. (2024) leverages LLMs to predict chunk boundaries more aligned with human interpretation. However, existing approaches often lack the flexibility to adapt to diverse query intents or varying deployment constraints. Moreover, some methods rely on computationally expensive large models, where the marginal gains of semantic chunking may not always justify the computational overhead Qu et al. (2025). Structurally, these approaches still operate within a single-granularity paradigm, consequently limiting their ability to simultaneously accommodate different information needs within the same document collection. Furthermore, the reliance on complex model inference for boundary detection introduces severe latency, rendering them impractical for large-scale real-time retrieval.
@@
To overcome the intrinsic constraints of fixed chunking schemes and the computational inefficiency of existing semantic approaches, FreeChunker is proposed in this paper—a novel Cross-Granularity Encoding Framework that introduces an alternative chunking paradigm. As illustrated in Figure 1, unlike traditional chunking methods that are typically constrained by fixed text boundaries, the proposed framework enables retrievers to extract arbitrary combinations of sentences from documents, thereby providing enhanced flexibility in content retrieval. The core contributions of this paper are summarized as follows:
@@
  • Computational efficiency: The framework treats sentences as atomic units, avoiding time-consuming boundary identification processes. Instead, it utilizes preset granularity combinations, significantly reducing both chunking and encoding time.
  • Superior performance: The framework achieves the best average retrieval performance by adapting multiple granularity levels to diverse query requirements, consistently outperforming existing baselines across various scenarios on the LongBench V2 benchmark.
@@ 2 Related Work
As a foundational step in RAG, text chunking substantially influences overall system performance. The most straightforward strategy divides documents into fixed-granularity chunks, which is simple to implement but often misaligns with semantic boundaries, consequently leading to information fragmentation or redundancy Ru et al. (2024). More critically, fixed chunking forces a single granularity choice across the entire document, thereby preventing the system from simultaneously accessing both fine-grained details and coarse-grained context that different queries may require. This granularity rigidity fundamentally limits the retrieval system’s ability to adapt to diverse information needs within a single document collection.
@@
To address the aforementioned semantic boundary issues, recent work has explored more adaptive chunking strategies. SemanticChunker LangChain (2024) relies on sentence-level embedding similarity to enhance intra-chunk semantic coherence by identifying natural breakpoints based on semantic similarity thresholds. Meta-Chunking Zhao et al. (2024) proposes two adaptive segmentation algorithms leveraging large language models’ logical awareness: Perplexity Chunking identifies logical boundaries by analyzing inter-sentence perplexity variations, while Margin Sampling Chunking determines boundaries by comparing model probability differences for segmentation decisions. LumberChunker Duarte et al. (2024) leverages large language models to determine chunk boundaries more aligned with human interpretation. However, despite their semantic awareness, these methods still operate under a single granularity paradigm—each document is ultimately represented at one chosen level of granularity, whether sentence-level, paragraph-level, or custom semantic units.
@@
Inspired by the concept of mixture of experts (MoE) Jacobs et al. (1991); Shazeer et al. (2017), recent methods like MoC Zhao et al. (2025) and MoG (Mix-of-Granularity) Zhong et al. (2025) attempt to enhance adaptability by treating different chunking strategies as expert modules and employing routing mechanisms to dynamically select the most suitable chunker based on query characteristics. However, these MoE-inspired methods essentially aggregate existing chunkers as experts; consequently, both the quality of semantic chunking and the degree of granularity control remain fundamentally bounded by the capabilities of the underlying chunker experts.
@@ 3 Methodology
In FreeChunker, sentences are treated as atomic units and chunk embeddings are generated across multiple granularities simultaneously. This approach enables retrieval systems to access any desired chunk size on demand, ranging from fine-grained single sentences to coarse-grained multi-sentence contexts.
Before detailing the mathematical formulation, a high-level intuition of the approach is provided. Unlike traditional methods that physically segment text and encode each chunk separately, FreeChunker employs a unified encoding process. Imagine a “mask” overlaid on the document’s sentence representations. By adjusting this mask, it is possible to dynamically specify which sentences should be attended to and combined into a chunk embedding. Crucially, by constructing a composite mask that contains patterns for multiple granularities (e.g., 1-sentence, 2-sentence, …, k-sentence chunks), embeddings for all desired chunk sizes can be generated in a single forward pass of the model. This eliminates the need for redundant computations and allows for flexible, overlapping chunk definitions without repeated encoding.
3.1 Sentenizer
The Sentenizer serves as the initial stage of the FreeChunker pipeline. Given a document $\mathcal{D}$, it first decomposes it into atomic sentence units $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$. Subsequently, a token-level embedding model $\mathcal{M}$ is employed to encode these sentences into a sentence embedding matrix $\mathcal{E}=[\vec{e_{1}},\vec{e_{2}},\dots,\vec{e_{n}}]\in\mathbb{R}^{n\times d}$, where $d$ denotes the embedding dimension. This process provides the foundational input for the subsequent cross-granularity encoding process.
@@ 3.2 Cross-Granularity Chunk Pattern
Based on the $n$ sentences obtained from the Sentenizer, a Chunk Pattern Mask $\mathbf{P}\in\{0,-\infty\}^{m\times n}$ is constructed, where $m$ is the number of desired chunks. Each row $\mathbf{P}[i,:]$ represents the $i$-th chunk pattern, and $\mathbf{P}[i,j]=0$ indicates that sentence $s_{j}$ belongs to that pattern.
@@
Figure 3: Training Dynamics and Generalization Across Embedding Models. Training and validation curves (top) illustrate training loss in blue and validation loss in orange, and test set loss distributions (bottom) show the cosine loss on an in-distribution test set, demonstrating consistent convergence and effective generalization.
Therefore, the framework supports the superposition processing of different granularities. As illustrated in Stage 2 of Figure 2, this results in a mask with a distinctive banded structure, where valid attention weights (0) are clustered around the diagonal for adjacent sentence chunks, while other areas are masked out (-$\infty$).
@@ 3.3 Cross-Granularity Encoder
After obtaining the Chunk Pattern Mask $\mathbf{P}$ and sentence embeddings $\mathcal{E}$, independently encoding each chunk would typically cause extensive redundant computation. For any chunk spanning from sentence $s_{i}$ to sentence $s_{j}$, its true embedding is conceptually equivalent to $\vec{e}_{i,j}=\mathcal{M}([s_{i},s_{i+1},\dots,s_{j}])$, where $s_{i},s_{i+1},\dots,s_{j}]$ represents the concatenation of sentences.
To efficiently compute these representations, a specialized transformer architecture is modified. For each layer, a learnable chunk embedding $\mathbf{h}_{\text{chk}}\in\mathbb{R}^{d}$ is introduced that is replicated $m$ times to form $\mathcal{H}=[\mathbf{h}_{\text{chk}},\mathbf{h}_{\text{chk}},\dots,\mathbf{h}_{\text{chk}}]^{T}\in\mathbb{R}^{m\times d}$. The cross-granularity attention is then computed using the sentence embeddings $\mathcal{E}$ as Keys and Values:
@@
Following the attention mechanism, standard Transformer layer operations are applied, including Layer Normalization and Feed-Forward Networks (FFN), to produce the final chunk embeddings. This design allows parallel generation of all chunk embeddings in a single forward pass, with sentence-level embeddings being reused across different granularities, thereby avoiding redundant encoding of sentences.
3.4 Deduplicate and Concatenate
To address the overlap in multi-granularity retrieved chunks, a post-processing strategy is adopted. Overlapping segments are deduplicated by marking the starting positions, and the retained unique segments are concatenated in order to reconstruct the narrative context.
@@ 4.1 Training Setup
To train the cross-granularity framework (Section 3) for effective chunk representation, three fundamental token-level embedding models are selected: jina-embeddings-v2-small-en (33M parameters), nomic-embed-text-v1.5 (137M parameters), and BGE-M3 (570M parameters). These models represent different scales and architectures, enabling a thorough assessment across varying embedding capacities.
@@
The training dataset comprises 100K documents sampled from The Pile Gao et al. (2020), each truncated to 8000 tokens and split into sentences. Training pairs $(\vec{e},\vec{v})$ are constructed across 5 granularities (2, 4, 8, 16, 32 sentences), where $\vec{e}$ is the ground-truth embedding obtained by encoding concatenated sentences using the base models, and $\vec{v}$ denotes the corresponding output from the proposed framework.